Alpha Mining Company recognizes $2 of depletion for each ton of ore mined. During 2016, 950,000 tons of ore were mined and 825,000 tons were sold. Using this information, how much should be recorded for FY 2016 for the following accounts:

1.accumulated Depletion

2.inventory

3.cost of Goods Sold
Answer #1

1. CALCULATE ACCUMLATED DEPLETION :

Accumlated depletion = 950000*2 = 1900000

2) CALCULATE INVENTORY

Inventory = (950000-825000)*2 = 250000

3) CALCULATE COST OF GOODS SOLD :

Cost of goods sold = 825000*2 = 1650000